Consolidated Industries is a diversified manufacturer with business units organized as divisions, including the Reigis Steel Division. Consolidated monitors its divisions on the basis of both unit contribution and return on investment (ROI), with investment defined as average operating assets employed. All investments in operating assets are expected to earn a minimum return of 12% before income taxes.

Reigis’s cost of goods sold is considered to be entirely variable; however, its administrative expenses do not depend on volume. Selling expenses are a mixed cost with one-third attributed to sales volume. The 2019 operating statement for Reigis follows. The division’s operating assets employed were $332,000,000 at November 30, 2019, unchanged from the year before.

REIGIS STEEL DIVISION
Operating Statement
For the Year Ended November 30, 2019
(000s omitted)
Sales revenue $ 186,000
Less expenses:
Cost of goods sold $ 96,000
Administrative expenses 19,000
Selling expenses 16,200 131,200
Income from operations, before tax $ 54,800

Required:

1. Calculate Reigis Steel Division’s unit contribution if it produced and sold 3,000,000 units during the year ended November 30, 2019. (Round your answer to 2 decimal places.)

2. Calculate the following performance measures for 2019 for Reigis:

a. Pretax ROI, based on average operating assets employed. (Round your answer to 2 decimal places.)

b. Residual income (RI), calculated on the basis of average operating assets employed. (Enter your answer in whole dollars, not in thousands.)
Answer #1

1.Total contribution margin = Sales Revenue – Variable costs

= 186,000-96000-16200*1/3

= $84,600 thousands

Unit contribution = Total contribution margin/Number of units

= 84600*1000/3,000,000

= $28.2 per unit

2.ROI = Income from operations before tax/Average Assets employed

= 54800*1000/332,000,000

= 16.5060%

i.e. 16.51%

3.Residual Income = Income from operations – Average Assets employed*Required rate of return

= 54800*1000 – 332,000,000*12%

= $14,960,000